OpenStudy (anonymous):

Evaluate y'. y=3^(x)*x^(3)

OpenStudy (anonymous):

\[y=3^xx^3\]

jimthompson5910 (jim_thompson5910):

use the product rule to get y=3^(x)*x^(3) y ' = d/dx[3^(x)*x^(3)] y ' = d/dx[3^(x)]*x^(3) + 3^(x)*d/dx[x^(3)] .. use product rule y ' = ln(3)*3^x*x^(3) + 3^(x)*3x^2 y ' = 3^(x)x^2 ( ln(3)*x + 3 ) ... this step is optional

OpenStudy (anonymous):

Thank you so much!!

jimthompson5910 (jim_thompson5910):

yw